OpenStudy (anonymous):

find f(x+h)-f(x)/(h) for the given function f. f(x)=9x+3

OpenStudy (anonymous):

it is the slope the slope of \(9x+3\) is ?

OpenStudy (anonymous):

9

OpenStudy (anonymous):

how do you know that the answer is the slope?

OpenStudy (kainui):

:)

OpenStudy (anonymous):

because this represents the change in \(y\) over the change in \(x\) which, for a line, is the slope we can do the algebra if you like

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

\[. f(x)=9x+3\] \[f(x+h)-f(x)=9(x+h)+3-(9x+3)=9x+9h+3-9x-3=9h\]

OpenStudy (anonymous):

\[\frac{f(x+h)-f(x)}{h}=\frac{9h}{h}=9\]